solve using elimination method: 3x-4y=9 and -3x+2y=9

Answer by mananth(16946)

3 x -4 y = 9 .............1
Total value
-3 x + 2 y = 9 .............2
add (1) & (2)
-2y=18
/2
y=-9
plug y in (1)
3x-4y=9
3x-4(-9)=9
3x+36=9
3x=-27
/3
x=-9